# Rigor in AI: Doing Rigorous AI Work Requires a Broader, Responsible AI-Informed Conception of Rigor

### v1

@@ Abstract
In AI research and practice, rigor remains largely understood in terms of methodological rigor—such as whether mathematical, statistical, or computational methods are correctly applied. We argue that this narrow conception of rigor has contributed to the concerns raised by the responsible AI community, including overblown claims about AI capabilities. Our position is that a broader conception of what rigorous AI research and practice should entail is needed. We believe such a conception—in addition to a more expansive understanding of (1) methodological rigor—should include aspects related to (2) what background knowledge informs what to work on (epistemic rigor); (3) how disciplinary, community, or personal norms, standards, or beliefs influence the work (normative rigor); (4) how clearly articulated the theoretical constructs under use are (conceptual rigor); (5) what is reported and how (reporting rigor); and (6) how well-supported the inferences from existing evidence are (interpretative rigor). In doing so, we also aim to provide useful language and a framework for much-needed dialogue about the AI community’s work by researchers, policymakers, journalists, and other stakeholders.
@@ 1 Rigor in AI Research and Practice
Rigor remains a subject of intense debate in science [e.g., 2, 41, 66, 101, 50, 126, 79, 104, 18], a debate we are unlikely to resolve here. We thus keep our goal relatively modest: to help broaden our collective perspective of what rigorous AI work should entail. We argue this is critically needed as relying on impoverished conceptions of rigor can have an undesirable, yet formative impact on the quality of both AI research and practice—heightening concerns ranging from unsubstantiated claims about AI systems [e.g., 124, 112, 138, 145, 33] to a plethora of unintended consequences [e.g., 123, 107, 78].
@@
Rigor in AI:  The debate surrounding rigor in science has by no means evaded the AI community [e.g., 129, 62, 85, 114]. In AI research and practice, rigor remains largely understood in terms of methodological rigor, which is typically conceptualized as whether mathematical, statistical, or computational methods are correctly applied; whether new methods, models, or systems are tested on large-scale or complex benchmarks and compared with a sufficient number of competing methods, models, or systems; whether the methods or analyses can scale or generalize; or whether the phenomena under analysis were—in contrast to more qualitative work—mathematically formalized and quantified [e.g., 57, 155, 139, 127, 24, 73, 47, 62, 129, 20]. These conceptualizations are often shaped by implicit assumptions that more complex methods and architectures and larger data samples are better [e.g., 7, 139, 136, 81, 45], by the common practice of relying on benchmark-driven evaluations [e.g., 155, 84, 141], and by a dominant mode of thinking oriented towards algorithmic formalism (centered on ideas of objectivity/neutrality, abstract and mathematical representations, and universalism/generalization) [e.g., 57, 20, 147, 87, 42]. Yet such conceptualizations of rigor may fail to demand, for instance, that benchmarks be fit-for-purpose or proven to measure what they claim to be measuring [e.g., 26, 110, 142], or that the knowledge or assumptions the work relies on be reliable or valid [e.g., 26, 112, 96, 90], putting into question the integrity and reliability of any conclusions drawn based on such benchmark-centered evaluations or that depend on questionable assumptions.
@@
While such failures threaten the scientific integrity of AI research, the consequences of putting insufficiently rigorous AI into practice are often seen as the purview of responsible AI—which is nevertheless generally seen as separate from rigor, as it is understood as more concerned with ethics, stakeholders, societal impacts and harms, and real-world deployment scenarios. However, it is often only in real-world deployment scenarios or when considering stakeholders that the inadequacies and impact of current approaches to (lack of) rigor in AI research and practice become clear. We thus recast this distinction, and argue that by making visible and demanding attention to such failures, responsible AI asks researchers and practitioners to uphold principles of scientific integrity in their work. Although responsible AI is often seen as out of scope for an AI researcher or practitioner not engaged in that space, any scientist should see rigor as well within scope. In other words, even though unrigorous AI work implicates what are often seen as responsible AI consequences, we argue that a broader notion of rigor that accounts for what produces these consequences needs to be considered by all AI researchers and practitioners.111 Positionality statement: In recent years, we have noticed increasing tensions within and across AI and AI-adjacent communities concerning what research questions should be prioritized, who can even be trusted to conduct certain types of work, and what rigorous research entails. We also observed trends in the types of work submitted and published at both AI and AI-adjacent venues, raising concerns about a pernicious lack of conceptual clarity, growing pockets of research motivated by hypothetical, speculative settings, and an increasing use of ambiguous, non-productive terminology. These experiences inform and motivate this paper. Our perspective is also informed by our diverse disciplinary backgrounds spanning computer systems, theoretical and applied ML, information retrieval, computational social science, natural language processing, computational linguistics, computer vision, human-computer interaction, law and policy, science and technology studies, and philosophy.
@@
We take the position that doing rigorous AI work requires broadening our understanding of what rigor in AI research and practice should entail by drawing on work by the responsible AI community. [^2] We use responsible AI to broadly refer to critical work on the impact of AI on people & society from several related communities, including ethical AI, responsible AI, ethics of AI, and science & technology studies. To this end, we foreground a wider range of critical considerations from responsible AI literature (broadly construed) for rigorous AI work. Instead of prescribing rigid criteria for rigorous AI work, our goal is to provide useful scaffolding for much-needed dialogue about and scrutiny of the community’s work by researchers as well as policymakers, journalists, and other stakeholders.
@@ 2 Facets of Rigor in AI Research and Practice
We foreground six key facets of rigor—epistemic, normative, conceptual, methodological, reporting, and interpretative—that we argue AI research and practice should consider. While it may be difficult to draw clear boundaries between some of these facets, we discuss them separately to provide distinct lenses for reflecting on and interrogating the quality and scientific integrity of AI work (see Table 1). Specifically, for each facet, we provide a descriptive overview—what the facet is concerned with, i.e., what the object of concern is for that facet—and an evaluative overview—what the facet asks for. All facets of rigor are inherently about the choices we make about an object of concern (e.g., epistemic rigor is concerned with background knowledge, while conceptual rigor is concerned with theoretical constructs). Having these distinct facets encourages researchers and practitioners to think carefully about the choices they make for each object of concern, and disentangles possible debates about these choices—e.g., separates debates about which norms we should follow (normative rigor) from what background knowledge should inform the work (epistemic rigor). For each facet, we discuss examples of mechanisms that help promote rigor along that facet, which can include a mix of processes and desiderata; while we foreground each mechanism for only one of the facets, we note that some mechanisms might help foster rigor along more than one facet (e.g., engaging with construct and internal validity concerns can foster both methodological and interpretative rigor).
@@
Figure 1 provides a simplified overview of the objects of concern for each facet of rigor and of possible common dependencies among them. It illustrates how limiting our conception of rigor to methodological concerns may obfuscate how our work and the claims we make are shaped by a variety of choices that both precede and succeed any methodological considerations, even when some of those choices we make remain tacit or implicit (e.g., as is often the case for common disciplinary norms, standards, or practices [155, 53]). The figure also underscores how it may be difficult to make good choices for “downstream” objects when “upstream” objects are poorly chosen. For instance, making poor construct choices (conceptual rigor) may reduce our chances of operationalizing those constructs well (methodological rigor). The different facets of rigor may, however, also be tangled in complex relations of mutual interdependency [e.g., 42, 54, 143]—as sometimes methodological choices may, for instance, in turn “limit the structure of one’s theoretical con[structs]” [42]—and present choices may (and usually do) impact future work. For example, a lack of interpretative rigor—when ambiguous or baseless claims are being made—can have epistemic consequences for future work relying on those claims [e.g., 63, 39]. We further unpack these below:
@@ 2.1 Epistemic rigor
Before considering any methodological questions, we have to contend with what it is that is being investigated, why it merits consideration, and what knowledge is “not under investigation, but [is] assumed, asserted, or essential” [58]. That is, we need to clarify the facts and background assumptions upon which the work relies in order to establish what the foundation for the work is and whether that foundation is sound. A frequently given example for epistemic failures is work about “the ability to predict unobservable latent character traits, including homosexuality, political ideology, and criminality, from photographs of human faces or other records of outward appearance” [8]. Such work assumes that we can predict actions (what one does) or inner character (what one likes, thinks, or values) based on physical attributes (how one looks). Although this assumption has its roots in “physiognomy”—regarded as pseudo-scientific as it relies on a set of epistemically baseless and extensively debunked claims [152, 134]—pockets of AI research recurrently draw upon it [8, 152]. This example also illustrates a broader class of epistemic failures: cases where a failure to scrutinize background assumptions may lead to work whose validity depends on whether existing methods, tools, or systems function—or can be made to function—on given tasks when they do not or cannot [112], including because the tasks are conceptually or practically impossible or because there is no reliable evidence that those methods, tools, or systems are fit-for-purpose or reliable [112, 155, 116, 96].
@@
Epistemic rigor is thus not only about making sure that the background knowledge that new work builds on and how that knowledge is acquired are appropriate and appropriately applied, but also that that knowledge is appropriately justified [34]. Rigorously applying statistical or computational methods (methodological rigor) will do little to mitigate epistemic concerns if the problems being tackled or the assumptions underpinning the use of some methods are baseless, nonsensical, unethical, or grounded in pseudo-scientific or scientifically shallow work [8, 112, 130, 78]. Why and what is being studied, built, or deployed [13]; why, what and whose problems are being prioritized [21, 121]; what implicit or tacit assumptions are being made about the stated problems or solutions under consideration [112, 91]; and whether the choices of problems and methods are drawing on valid and well-founded evidence [22] are all examples of the types of key epistemic considerations that we should engage with. By contrast, easily available or common artifacts (such as datasets, methods, tools, or systems) tend to serve as “tools of opportunity, not instruments of epistemic rigor” [151], with researchers and practitioners often prioritizing work based on whether there are some existing artifacts available without appropriately scrutinizing the knowledge and assumptions underpinning them [155].
@@
Epistemic rigor, however, does not necessarily require specific epistemological commitments or choices but rather that those commitments and choices be made explicit. While doing so may not definitively answer whether some problems or assumptions are baseless, nonsensical, or unethical, articulating epistemic commitments and the existing knowledge the work relies on lays down the grounds on which people can have discussions and work out disagreements [30].
@@
Mechanisms to promote epistemic rigor Ensuring work is appropriately grounded in past literature [e.g., 8] and that underlying assumptions are made explicit and appropriately interrogated [e.g., 112, 155, 91] can help foster epistemic rigor, and in turn all other facets of rigor (§2.2–§2.6).
@@
Appropriate grounding:  Common goals in AI often include producing new insights, theories, or artifacts, providing evidence in support of existing theories, evaluating existing artifacts, or debunking prior work. A failure to review, appropriately situate within, and acknowledge prior literature, and how it influenced the questions being asked or the solutions being considered, can cast doubt on whether any meaningful progress towards those goals was actually made [e.g., 60, 66]. When such failures become systematic, research communities risk “curl[ing] up upon themselves and become self-referential systems that orient more [internally]” [35] and developing their separate “terminology, source texts, and knowledge claims” [5]. Indeed, concerns that “[f]indings within [a research] community are often self-referential and lack [quality]” [65] are critical as this can make it difficult to trace the provenance of central claims that form the foundation that new work builds on or is motivated by. It can also result in an overall narrowing of what questions AI research tackles and the epistemic marginalization of certain viewpoints [80, 4, 96]. Appropriate grounding requires due diligence when reviewing and acknowledging work in one’s subfield and related fields [16, 66, 99].
@@
Interrogating assumptions:  Any research work necessarily relies on assumptions about what is important, what is possible, or what represents sufficient or useful evidence [e.g., 115]. Not only does a lack of epistemic rigor make it hard to situate new work and the knowledge it produces in the context of existing knowledge it may draw from, relate to, corroborate, or dispute, but a lack of due diligence about the background knowledge underlying the work further risks bringing to the fore long-debunked claims from other disciplines—e.g., risking the reanimation of physiognomic methods [8, 134, 152]. Interrogating the assumptions underpinning such tasks—e.g., why would outer attributes be useful proxies for someone’s character?—can expose them as a category error where observable traits (e.g., skin tone) are incorrectly treated as direct indicators of internal states (e.g., personality). Making background assumptions explicit thus facilitates our ability to scrutinize them [155, 12].

### v2

@@ Abstract
In AI research and practice, rigor remains largely understood in terms of methodological rigor—such as whether mathematical, statistical, or computational methods are correctly applied. We argue that this narrow conception of rigor has contributed to the concerns raised by the responsible AI community, including overblown claims about the capabilities of AI systems. Our position is that a broader conception of what rigorous AI research and practice should entail is needed. We believe such a conception—in addition to a more expansive understanding of (1) methodological rigor—should include aspects related to (2) what background knowledge informs what to work on (epistemic rigor); (3) how disciplinary, community, or personal norms, standards, or beliefs influence the work (normative rigor); (4) how clearly articulated the theoretical constructs under use are (conceptual rigor); (5) what is reported and how (reporting rigor); and (6) how well-supported the inferences from existing evidence are (interpretative rigor). In doing so, we also provide useful language and a framework for much-needed dialogue about the AI community’s work by researchers, policymakers, journalists, and other stakeholders.
@@ 1 Rigor in AI Research and Practice
Rigor remains a subject of intense debate in science [e.g., adler2012disciplining, coryn2007holy, houston2019four, national2019reproducibility, forero2018application, schaller2016empirical, kilicoglu2018biomedical, nowell2019reviewer, bernieri1991rigor], a debate we are unlikely to resolve here. We thus keep our goal relatively modest: to help broaden the AI community’s perspective of what rigorous AI [^1] Throughout this paper, we deliberately leave the term AI under-specified in order to capture the full constellation of meanings the term has been used for, and thus any work that practitioners or researchers would describe as AI—i.e., if they consider themselves as working on AI, our call is for them. We thus mainly use the term as a modifier to help scope the community this call is addressing, and the work and artifacts this community does or uses. work should entail. We argue this is critically needed as relying on impoverished conceptions of rigor can have an undesirable, yet formative impact on the quality of both AI research and practice—heightening concerns ranging from unsubstantiated claims about AI systems [e.g., schaeffer2023emergent, raji2022fallacy, van2024reclaiming, wang2024apo, buyl2025ai] to a plethora of unintended consequences [e.g., saxena2025ai, olteanu2023responsible, keyes2019mulching].
@@
Rigor in AI:  The debate surrounding rigor in science has by no means evaded the AI community [e.g., sculley2018winner, herrmann2024position, liao2021we, recht2023]. In AI research and practice, rigor remains largely understood in terms of methodological rigor, which is typically conceptualized as whether mathematical, statistical, or computational methods are correctly applied; whether new methods, models, or systems are tested on large-scale or complex benchmarks and compared with a sufficient number of competing methods, models, or systems; whether the methods or analyses can scale or generalize; or whether the phenomena under analysis were—in contrast to more qualitative work—mathematically formalized and quantified [e.g., green2020algorithmic, zhou2022deconstructing, varoquaux2024hype, Sculley2023SurveyOS, blili2025stop, kaplan2020scaling, dziugaite2020search, herrmann2024position, sculley2018winner, birhane2022values]. These conceptualizations are often shaped by implicit assumptions that more complex methods and architectures and larger data samples are better [e.g., anderson2008end, varoquaux2024hype, sutton2019bitter, krizhevsky2012imagenet, diaz2024scaling], by the common practice of relying on benchmark-driven evaluations [e.g., zhou2022deconstructing, liangholistic, wagstaff2012machine], and by a dominant mode of thinking oriented towards algorithmic formalism (centered on ideas of objectivity/neutrality, abstract and mathematical representations, and universalism/generalization) [e.g., green2020algorithmic, birhane2022values, weerts2024neutrality, lipton2019troubling, danziger1985methodological]. Yet such conceptualizations of rigor may fail to demand, for instance, that benchmarks be fit-for-purpose or proven to measure what they claim to be measuring [e.g., blodgett2021stereotyping, porada2024challenges, wallach2025position], or that the knowledge or assumptions the work relies on be reliable or valid [e.g., blodgett2021stereotyping, raji2022fallacy, messeri2024artificial, lu2022subverting]. This puts into question the integrity and reliability of any conclusions drawn based on such benchmark-centered evaluations or that depend on questionable assumptions.
@@
While such failures threaten the scientific integrity of AI research, foreseeing and addressing the consequences of putting insufficiently rigorous AI artifacts (e.g., models, systems, applications, outputs, data) into practice are often seen as the purview of responsible AI—which is nevertheless generally seen as separate from rigor, as it is understood as more concerned with ethics, stakeholders, societal impacts and harms, and real-world deployment scenarios. However, it is often only in real-world deployment scenarios or when considering stakeholders that the inadequacies and impact of current approaches to (lack of) rigor in AI research and practice become clear. We thus recast this distinction, and argue that by making visible and demanding attention to such failures, responsible AI asks researchers and practitioners to uphold principles of scientific integrity in their work. Although responsible AI is often seen as out of scope for an AI researcher or practitioner not engaged in that space, any scientist should see rigor as well within scope. In other words, even though unrigorous AI work implicates what are often seen as responsible AI consequences, we argue that a broader notion of rigor that accounts for what produces these consequences needs to be considered by all AI researchers and practitioners. [^2] Positionality statement: In recent years, we have noticed increasing tensions within and across AI and AI-adjacent communities concerning what research questions should be prioritized, who can even be trusted to conduct certain types of work, and what rigorous research entails. We also observed trends in the types of work submitted and published at both AI and AI-adjacent venues, raising concerns about a pernicious lack of conceptual clarity, growing pockets of research motivated by hypothetical, speculative settings, and an increasing use of ambiguous, non-productive terminology. These experiences inform and motivate this paper. Our perspective is also informed by our diverse disciplinary backgrounds spanning computer systems, theoretical and applied ML, information retrieval, computational social science, natural language processing, computational linguistics, computer vision, human-computer interaction, law and policy, science and technology studies, and philosophy.
@@
We take the position that doing rigorous AI work requires broadening our understanding of what rigor in AI research and practice should entail by drawing on work by the responsible AI community. [^3] We use responsible AI to broadly refer to critical work on the impact of AI artifacts on people and society from several communities, including ethical AI, responsible AI, ethics of AI, and science and technology studies. To this end, we foreground a wider range of critical considerations from responsible AI literature (broadly construed) for rigorous AI work. Instead of prescribing rigid criteria for rigorous AI work, our goal is to provide useful scaffolding for much-needed dialogue about and scrutiny of the community’s work by researchers as well as policymakers, journalists, and other stakeholders.
@@ 2 Facets of Rigor in AI Research and Practice
We foreground six key facets of rigor—epistemic, normative, conceptual, methodological, reporting, and interpretative—that we argue AI research and practice should contend with. While it may be difficult to draw clear boundaries between some of these facets, we discuss them separately to provide distinct lenses for reflecting on and interrogating the quality and scientific integrity of AI work (see Table 1). Specifically, for each facet, we provide a descriptive overview—what the facet is concerned with, i.e., what the object of concern is for that facet—and an evaluative overview—what the facet asks for. All facets of rigor are inherently about the choices we make about an object of concern (e.g., epistemic rigor is concerned with background knowledge, while conceptual rigor is concerned with theoretical constructs). Having these distinct facets encourages researchers and practitioners to think carefully about the choices they make for each object of concern, and disentangles possible debates about these choices—e.g., it separates debates about which norms we should follow (normative rigor) from what background knowledge should inform the work (epistemic rigor). For each facet, we discuss examples of mechanisms that help promote rigor along that facet, which can include a mix of processes and desiderata. While we foreground each mechanism for only one of the facets, we note that some mechanisms might help foster rigor along more than one facet (e.g., engaging with construct and internal validity concerns can foster both methodological and interpretative rigor).
@@
Figure 1 provides a simplified overview of the objects of concern for each facet of rigor and of possible common dependencies among them. It illustrates how limiting our conception of rigor to methodological concerns may obfuscate how our work and the claims we make are shaped by a variety of choices that both precede and succeed any methodological considerations, even when some of those choices remain tacit or implicit (e.g., as is often the case for common disciplinary norms, standards, or practices [zhou2022deconstructing, geiger2021garbage]). The figure also underscores how it may be difficult to make good choices for “downstream” objects when “upstream” objects are poorly chosen. For instance, making poor construct choices (conceptual rigor) may reduce our chances of operationalizing those constructs well (methodological rigor). The different facets of rigor may, however, also be tangled in complex relations of mutual interdependency [e.g., danziger1985methodological, george1973conceptualization, wallach2025position]—as methodological choices may, for instance, in turn “limit the structure of one’s theoretical con[structs]” [danziger1985methodological] if the methods determine what is observed and thus theorized about—and present choices may (and usually do) impact future work. For example, a lack of interpretative rigor—when ambiguous or baseless claims are being made—can have epistemic consequences for future work relying on those claims [e.g., herrmann2024position, cooper2021hyperparameter]. We further unpack these below:
@@ 2.1 Epistemic rigor
Before considering any methodological questions, we have to contend with what it is that is being investigated, why it merits consideration, and what knowledge is “not under investigation, but [is] assumed, asserted, or essential” [guest2024makes]. That is, we need to clarify the facts and background assumptions upon which the work relies in order to establish what the foundation for the work is and whether that foundation is sound. A frequently given example for epistemic failures is work asserting that records of physical appearance (e.g., photographs of faces) can be used to predict latent character traits (e.g., sexuality, political ideology, or criminality) [andrews2024reanimation]. Such work assumes that actions (what one does) or inner character (what one likes, thinks, or values) can be predicted based on physical attributes (how one looks). Although this assumption has its roots in “physiognomy”—regarded as pseudo-scientific as it relies on a set of epistemically baseless and extensively debunked claims [y2023physiognomy, stark2021physiognomic]—pockets of AI research recurrently draw upon it [andrews2024reanimation, y2023physiognomy]. This example also illustrates a broader class of epistemic failures: cases where a failure to scrutinize background assumptions may lead to work whose validity depends on whether existing methods, tools, or systems function—or can be made to function—on given tasks when they do not or cannot [raji2022fallacy], including because the tasks are conceptually or practically impossible or because there is no reliable evidence that those methods, tools, or systems are fit-for-purpose or reliable [raji2022fallacy, zhou2022deconstructing, roberts2021common, messeri2024artificial].
@@
Epistemic rigor is thus not only about making sure that the background knowledge new work builds on and how that knowledge is acquired are appropriate and appropriately applied, but also that that knowledge is appropriately justified [carter2007justifying]. Rigorously applying statistical or computational methods (methodological rigor) will do little to mitigate epistemic concerns if the problems being tackled or the assumptions underpinning the use of some methods are baseless, theoretically implausible, nonsensical, or grounded in pseudo-scientific or scientifically shallow work [andrews2024reanimation, raji2022fallacy, sloane2022silicon, keyes2019mulching, adolfi2024empirical]. Why and what is being studied, built, or deployed [barocas2020not]; why, what and whose problems are being prioritized [birhane2022forgotten, sambasivan2021everyone]; what implicit or tacit assumptions are being made about the stated problems or solutions under consideration [raji2022fallacy, lucy2024one, adolfi2024empirical]; and whether the choices of problems and methods are drawing on valid and well-founded evidence [birhane2023science] are all examples of the types of key epistemic considerations that we should engage with. By contrast, easily available or common artifacts (such as datasets, methods, tools, or systems) tend to serve as “tools of opportunity, not instruments of epistemic rigor” [winsberg2017success], with researchers and practitioners often prioritizing work based on whether there are some existing artifacts available without appropriately scrutinizing the knowledge and assumptions underpinning them [zhou2022deconstructing].
@@
Epistemic rigor, however, does not necessarily require specific epistemological commitments or choices but rather that those commitments and choices be made explicit. While doing so may not definitively answer whether some problems or assumptions are baseless, nonsensical, or unethical, articulating epistemic commitments and the existing knowledge the work relies on lays down the grounds on which people can have discussions and work out disagreements [brister2016disciplinary].
@@
Mechanisms to promote epistemic rigor Ensuring work is appropriately grounded in past literature [e.g., andrews2024reanimation] and that underlying assumptions are made explicit and appropriately interrogated [e.g., raji2022fallacy, zhou2022deconstructing, lucy2024one] can help foster epistemic rigor, and in turn all other facets of rigor (§2.2–§2.6).
@@
Appropriate grounding:  Common goals in AI research often include producing new insights, theories, or artifacts, providing evidence in support of existing theories, evaluating existing artifacts, or debunking prior work. A failure to review, appropriately situate within, and acknowledge prior literature, and how it influenced the questions being asked or the solutions being considered, can cast doubt on whether any meaningful progress towards those goals was actually made [e.g., Hadfield-Menell2024need, houston2019four]. When such failures become systematic, research communities risk “curl[ing] up upon themselves and becom[ing] self-referential systems that orient more [internally]” [cetina2007culture] and developing their separate “terminology, source texts, and knowledge claims” [ahmed2024field]. Indeed, concerns that “[f]indings within [a research] community are often self-referential and lack [quality]” [sciencecommittee2023] are critical as this can make it difficult to trace the provenance of central claims that form the foundation that new work builds on or is motivated by. It can also result in an overall narrowing of what questions AI research tackles and the epistemic marginalization of certain viewpoints [klinger2020narrowing, ahmed2024narrow, messeri2024artificial]. Appropriate grounding requires due diligence when reviewing and acknowledging work in one’s subfield and related fields [bender2021dangers, houston2019four, mitchell2024oversight].
@@
Interrogating assumptions:  Any research work necessarily relies on assumptions about what is important, what is possible, or what represents sufficient or useful evidence [e.g., rehak2021language]. Not only does a lack of epistemic rigor make it hard to situate new work and the knowledge it produces in the context of existing knowledge it may draw from, relate to, corroborate, or dispute, but a lack of due diligence about the background knowledge underlying the work further risks bringing to the fore long-debunked claims from other disciplines—e.g., risking the reanimation of physiognomic methods [andrews2024reanimation, stark2021physiognomic, y2023physiognomy]. Interrogating the assumptions underpinning such tasks—e.g., why would outer attributes be useful proxies for someone’s character?—can expose them as a category error where observable traits (e.g., skin tone) are incorrectly treated as direct indicators of internal states (e.g., personality). Making background assumptions explicit thus facilitates our ability to scrutinize them [zhou2022deconstructing, balayn2023faulty].